### v1

# Modeling Multimodal Social Interactions: New Challenges and Baselines with Densely Aligned Representations

Abstract
Understanding social interactions involving both verbal and non-verbal cues is essential to effectively interpret social situations. However, most prior works on multimodal social cues focus predominantly on single-person behaviors or rely on holistic visual representations that are not densely aligned to utterances in multi-party environments. They are limited in modeling the intricate dynamics of multi-party interactions. In this paper, we introduce three new challenging tasks to model the fine-grained dynamics between multiple people: speaking target identification, pronoun coreference resolution, and mentioned player prediction. We contribute extensive data annotations to curate these new challenges in social deduction game settings. Furthermore, we propose a novel multimodal baseline that leverages densely aligned language-visual representations by synchronizing visual features with their corresponding utterances. This facilitates concurrently capturing verbal and non-verbal cues pertinent to social reasoning. Experiments demonstrate the effectiveness of the proposed approach with densely aligned multimodal representations in modeling social interactions. We will release our benchmarks and source code to facilitate further research.
1 Introduction
Real-world social interactions involve intricate behaviors between multiple people. People communicate not only through spoken language, but also through non-verbal cues like gestures. While spoken language conveys direct meaning, inferring social meaning from it alone is sometimes vague. Non-verbal cues often play a crucial role in clarifying subtle social nuances. Therefore, comprehensively understanding social interactions involving multimodal social cues is essential to interpret social situations appropriately.
Social deduction games, where players take on roles and try to reveal the hidden roles of their opponents, provide an effective testbed for studying multimodal social interactions. These games require players to engage in communication, deception, inference, and collaboration, encompassing rich social interactions. These social interactions comprise various multimodal cues including verbal (e.g., language) and non-verbal (e.g., gesture, gaze) communications. Modeling such multimodal interactions in social deduction games can facilitate developing social artificial intelligence that can engage naturally alongside humans.
There have been attempts to investigate social behaviors in multimodal aspects by considering language and visual cues together. Some works tried to learn the relationships between spoken language and visual gestures for gesture generation [1, 32, 2] and gesture-language grounding [29]. Other multimodal approaches utilized the interconnection between spoken language and visual expressions to recognize human sentiment and emotion [53, 42, 22, 39, 21, 9]. However, these works focus mainly on the behavior of a single person or rely on holistic visual representations in multi-party environments, rather than modeling the fine-grained dynamics of social interactions among multiple people.
Recently, a multimodal work [28] addressed social behaviors in social deduction games where multiple players actively interact with one another. This work leveraged language and visual cues to predict persuasion strategies at the utterance level. However, it has limitations in modeling social interactions in terms of its task and methodology. Regarding the task, although persuasion strategies emerge in communication, it is largely a matter of understanding the social behavior of a single person rather than the dynamics among multiple people. In addition, it is difficult for their approach to distinguish and recognize fine-grained interactions because it utilizes holistic visual representations for the whole scene, despite multiple players being present.
To address these issues, we introduce three new multimodal tasks that encompass the dynamics of player interactions, along with a corresponding baseline model. We extend the social deduction game datasets [28] with additional extensive data annotations to curate new social tasks. Our curated tasks focus on identifying referents in multi-party conversations, a critical aspect of understanding social interactions. Figure 1 shows the overall concepts of our curated tasks. The three curated tasks are as follows.
1.
Speaking target identification: Identifying who a speaker is talking to in a conversation.
2.
Pronoun coreference resolution: Determining who a pronoun refers to in a conversation.
3.
Mentioned player prediction: Predicting who is mentioned by a name in a conversation.
These tasks are challenging as they require understanding the fine-grained dynamics of player interactions. For instance, when an utterance is accompanied by visual cues such as pointing gestures, it is required to figure out who is the speaker and who is being pointed at within the visual scene based on the context of utterances. This entails matching visually identified individuals with their corresponding references in spoken utterances, which enables understanding verbal and non-verbal cues comprehensively. Therefore, it is needed to align utterances with the corresponding player visuals and to utilize such densely aligned multimodal representations to tackle the tasks. Figure 2 shows an example of densely aligned representations.
To this end, we propose a novel baseline model leveraging densely aligned multimodal representations to address these new social tasks. We detect and visually track each player appearing in the video frames to distinguish individual behaviors. By establishing an initial alignment between player visual positions and their utterances, we continuously track player visuals in sync with dialogues. This alignment allows us to visually identify the speaker and the other players from given utterances. Encoding the visual gestures of a speaker and the relative positions of the other players allows us to decipher visual spatial relationships crucial for understanding non-verbal cues. Consequently, we can predict intended referents effectively by leveraging densely aligned multimodal representations from utterance language features and the corresponding visual features.
The major contributions of this paper are as follows.
We introduce new tasks in social deduction games via extensive data annotations: speaking target identification, pronoun coreference resolution, and mentioned player prediction. These tasks are challenging as they require understanding the fine-grained dynamics of interactions.
We propose a novel multimodal baseline for understanding social interactions in environments where multiple people actively interact with each other. To the best of our knowledge, this is the first work to address the dense alignment issue between language and visual social cues.
| Annotation Type | Utterance Example | YouTube DB | Ego4D DB | ||
| Count | Krippendorff’s | Count | Krippendorff’s | ||
| Speaking Target Identification | Why are you helping the Werewolves out? (To [Name]) | 3,255 | 0.922 | 832 | 0.907 |
| Pronoun Coreference Resolution | I’m a Villager which makes me think he ( [Name]) was the Werewolf | 2,679 | 0.962 | 503 | 0.846 |
| Mentioned Player Prediction | ”I’m the troublemaker and I switched [Name] with somebody” | 3,360 | – | 472 | – |
2 Related Work
2.1 Social Behavior Analysis
Analyzing social behaviors has been widely investigated in the fields of computer vision and natural language processing. Various works have focused primarily on analyzing social behaviors from a single-modal perspective. In terms of visual cues, some works proposed gaze target estimation techniques [11, 18, 27, 48, 47] to analyze where a person is looking within a scene. There have also been studies that recognize social gaze patterns between multiple people such as identifying shared attention [20, 16, 45, 36]. Gesture recognition approaches [54, 31, 56, 30, 3] have been researched to identify specific types of human gestures such as shaking hands and thumbs-up. Regarding language cues, dialogue act recognition methods [46, 50, 40, 41, 8, 34] have been introduced to understand the communicative intent behind utterances in social dialogues. Furthermore, there have been works for sentiment analysis and emotion recognition based on dialogue language [23, 4, 44, 57, 55].
Recently, joint modeling of visual and language modalities has been studied for social behavior analysis. Some works focused on learning the relationships between spoken languages and gestures for gesture generation [1, 32, 2] and gesture-language grounding [29]. Liu et al. [32] proposed a multimodal model that integrates visual, language, and speech cues via hierarchical manners to synthesize naturalistic gestures. Additionally, the intersection of spoken utterances and visual expressions has been explored for sentiment analysis and emotion recognition [53, 42, 22, 39, 21, 9]. Hu et al. [22] proposed a unified feature space to capture the knowledge of sentiment and emotion comprehensively from multimodal cues. There also have been multimodal works for question & answering in social contexts [52, 51, 37].
However, these works mainly focus on the behaviors of a single person or rely on holistic visual features that are not densely aligned to languages in multi-party environments. They are unable to model the complex dynamics of interactions, which requires understanding the spatial relationships of multiple people in addition to their utterances. We propose a novel baseline leveraging densely aligned language-visual representations to capture the fine-grained dynamics.
2.2 Social Deduction Game Modeling
There have been works to investigate computational models for social deduction games where players actively communicate and strategize with one another. Some prior studies have focused on developing game-playing agents and analyzing optimal strategies using game theory [6, 35, 5, 43, 12]. These works aim to model the state of the game computationally but do not address understanding the dialog and behaviors of players. Chittaranjan et al. [10] modeled game outcomes from communication patterns such as player speaking and interrupting behaviors. Bakhtin et al. [14] built an agent that can play diplomacy game by utilizing language models with strategic reasoning. These approaches do not capture verbal and non-verbal multimodal aspects of modeling social behaviors. Recently, Lai et al. [28] addressed social behaviors in social deduction games using multimodal representations. They leveraged language and visual cues to predict persuasion strategies at the utterance level such as identity declaration and interrogation.
However, this multimodal work is limited in addressing multi-person dynamics due to the lack of recognizing person-level features. To address this gap, we introduce three new benchmark tasks that explicitly demand recognizing the fine-grained dynamics between individual players across both language and visual representations.
3 Proposed Benchmark
3.1 Base Datasets
We extend two social deduction game datasets [28]: YouTube and Ego4D with additional extensive data annotations for curating new social tasks.
YouTube dataset. This dataset was collected from the YouTube video platform by searching keywords of Werewolf social deduction game. It contains 151 games of One Night Ultimate Werewolf, which corresponds to 151 separate videos with 14.8 hours. It consists of videos, transcripts, player roles, voting outcomes, and persuasion strategy labels. The transcripts comprise 20,832 utterances.
Ego4D dataset. This Ego4D dataset is a subset of Ego4D Social dataset [19]. It has 40 games of One Night Ultimate Werewolf and 8 games of The Resistance: Avalon. It contains 101 separate videos with 7.3 hours. Among them, we leverage 83 videos where we can visually identify individuals for new data annotations. To guarantee the visibility of all players within the frame, this dataset adopts third-person view videos instead of first-person view videos. It also consists of videos, transcripts, player roles, voting outcomes, and persuasion strategy labels. The transcripts contain 5,815 utterances during the game.
3.2 Data Annotation
To address the fine-grained dynamics of interactions, we design three new tasks in social deduction games: speaking target identification, pronoun coreference resolution, and mentioned player prediction. Annotators reference both transcripts and videos comprehensively to make their annotations in the transcripts. To achieve reliable quality of annotations, we initially request three annotators to label subsets of data and measure their annotation agreement using Krippendorff’s alpha [26]. After we train the annotators sufficiently with larger than 0.8, we request the three annotators to label independently for the rest of the data. Note that Krippendorff’s alpha 0.8 is generally considered to have good reliability with a high level of agreement [7].
Table 1 shows the summary of our data annotation results. We achieve sufficiently high for both speaking target identification and pronoun coreference resolution. Note that we obtain the annotations for the mentioned player prediction automatically. In the training and testing process, we anonymized all names in transcripts for our tasks (e.g., [“David”, “Alice”, “Thomas”] [“Player1”, “Player2”, “Player3”]). For each task, a test set is constructed using about 20% of the annotations. We split the training and test sets at the video level rather than at the individual utterance level to ensure no overlap in terms of contextual utterances. Annotation details for each task are as follows.
Task1: Speaking target identification. This task is to predict who a speaker is talking to in a dialogue. To this end, we annotate the labels of who a speaker is speaking to at the utterance level. Since utterances are often directed to all players, we performed the annotation on only the utterances that include [“you”, “your”]. We give our annotators three labeling options: (To Player#), (To Everyone), and (To Unknown). Based on the annotations, we added “(To [MASK])” at the end of a target utterance.
Task2: Pronoun coreference resolution. This task is to predict who a pronoun refers to in a dialogue. We conduct the annotation on the third-person pronouns that are used in our dataset, which are [“he”, “she”, “him”, “her”, “his”] in utterances. We give two options to annotators: [Player#] and [Unknown]. We only target the pronouns that indicate a specific player in the game. In terms of modeling the task, we changed a pronoun into [MASK] in a target utterance.
Task3: Mentioned player prediction. This task is to predict who is referred to by their name in a dialogue. Since we know the ground truth names in utterances, it is possible to annotate these mentioned player labels automatically. We mask a mentioned player name with [MASK] in utterances and predict who is referred to in that part.
4 Proposed Approach
The introduced social tasks can be formulated as follows. Let denote utterance texts that include the -th target utterance with a [MASK] token representing an unknown player, while indicates the corresponding video frames aligned with the utterance timeline. Given and , our objective is to optimize the multimodal predictive function to effectively classify the identity of the player associated with the [MASK].
To this end, we introduce a novel multimodal baseline that leverages densely aligned representations between language and visual cues to capture the fine-grained dynamics of interactions. Figure 3 shows the overall framework of the proposed multimodal baseline. The proposed model consists of four main parts: (i) language-visual alignment, (ii) visual interaction modeling, (iii) conversation context modeling, and (iv) aligned multimodal fusion for prediction.
4.1 Language-Visual Alignment
A key challenge in multimodal analysis for social interactions is establishing fine-grained alignments between visual and language cues. To address this, we introduce a technique to obtain densely aligned multimodal representations from utterances and the corresponding visual cues.
We detect and track players visually in video frames overtime using AlphaPose framework [17]. Once we initially match player visuals with the player references in the utterances, we can continuously identify players in both visual and language domains. This allows us to match who is speaking in the transcript with the location of the speaker in the video, and to understand where the listeners are located relative to the speaker. By employing language-visual alignment, we can achieve densely aligned representations that integrate both verbal and non-verbal cues. It enables us to tackle our social tasks effectively, allowing for a more nuanced and holistic understanding of interactions.
4.2 Visual Interaction Modeling
To distinguish individual players on video frames , we use the human pose keypoints from AlphaPose. Specifically, we extract 17 body keypoints (, ) for each player. Figure 3 shows the procedure of encoding visual interactions. The upper path (green) of Figure 3 indicates encoding a kinesics feature of a speaker while the middle path (purple) represents encoding spatial positions of players.
First, we use the keypoints of a speaker in the upper path of Figure 3. Among 17 keypoints, we leverage [nose, l-eye, r-eye, l-shoulder, r-shoulder, l-elbow, r-elbow, l-wrist, r-wrist] which are closely related to gaze and gesture characteristics. Let , denote the image coordinates of a part at time . For example, , indicates the nose point. To represent human motion in a unified coordinate, we normalize speaker keypoints by subtracting the speaker nose point from each part point. Each point vector , is independently encoded by an MLP point encoder into part point feature ( is channel dim). These part point features are concatenated and processed by an MLP kinesics encoder to obtain a speaker kinesics feature as follows.
Since we have multiple time steps, we can obtain .
In the meantime, the middle path (purple) of Figure 3 receives the position of each player. We consider the nose point of each player as their representative position. We normalize their nose points by subtracting the speaker’s nose point from them to get their relative positions from the speaker. We utilize their representative positions at a single time step corresponding to the start of the utterance. Let , denote the representative position of player #. Each point vector , is independently fed to an MLP point encoder to get a player point feature . We concatenate the player point features and feed them to an MLP position encoder to get . We then make aware of speaker knowledge. To this end, we obtain a speaker-label feature by passing a speaker-label one-hot vector through an FC layer. We combine with to obtain player position feature . These procedures are formulated as follows.
where indicates the maximum player number in the datasets (). If the number of players is less than for the current input data, we apply zero padding to the excess. If a player is temporarily undetected (e.g., offscreen for a short time), we proceed with position encoding by substituting the corresponding player position stored in a buffer to correct the player position.
Based on the speaker kinesics features and player position feature , we encode the visual interaction by capturing speaker kinesics motion with the context of player visual positions. and are passed through a visual interaction encoder sequentially which has the form of the transformer [49]. allows modeling dependencies between the speaker kinesics and player positions across time via self-attention. Finally, we can obtain visual interaction features that represent dynamics between the speaker and players based on the kinesics and their positions.
4.3 Conversation Context Modeling
The lower path (red) of Figure 3 shows encoding spoken utterances from players. To incorporate conversation context, we use surrounding utterances including the target utterance. The input to the language path is formulated as.
where denotes the target -th utterance, and the others indicate the preceding and following utterances. Note that the target utterance is the one that contains [MASK]. A [CLS] token is inserted in front of while a [SEP] token is inserted at the end of each utterance in for language tokenization processing. Note that all player names in utterances are anonymized as “[Player#]”. We leverage pre-trained language models based on masked-language modeling such as BERT [24]. The tokenized sequence of the utterances is fed into the language model. The output feature corresponding to the index of the [MASK] token is then retrieved. After passing it through an FC layer to match the channel dimension of the visual interaction features, we get a conversation context feature which contains the context around the [MASK].
| Method | Densely Aligned? | Speaking Target Identification (%) | |
| YouTube | Ego4D | ||
| BERT [24] | - | 65.8 | 56.8 |
| BERT + DINOv2 [38] | ✗ | 66.4 | 58.0 |
| BERT + MViT (Lai et al. [28]) | ✗ | 66.9 | 57.4 |
[…]
| RoBERTa [33] | - | 72.4 | 63.6 |
| RoBERTa + DINOv2 [38] | ✗ | 72.7 | 62.5 |
| RoBERTa + MViT (Lai et al. [28]) | ✗ | 73.1 | 64.2 |
[…]
| ELECTRA [13] | - | 65.8 | 60.8 |
| ELECTRA + DINOv2 [38] | ✗ | 65.3 | 60.2 |
| ELECTRA + MViT (Lai et al. [28]) | ✗ | 64.6 | 60.8 |
[…]
4.4 Aligned Multimodal Fusion
To fuse the aligned visual interaction features and conversation context feature , we first concatenate them along the sequence dimension along with a [AGG] token for feature aggregation. It can be formulated as follows.
Note that positional encoding [49] for transformers is applied to parts. Then, is processed with a multimodal transformer to encode their joint relationships. We leverage an output multimodal feature from the transformed [AGG] token. Finally, a densely aligned multimodal feature is passed through a classification head consisting of an FC layer and softmax to predict the anonymized player identity (e.g., Player#) for the target [MASK]. We optimize the model using cross-entropy loss between the predicted player and ground-truth label .
At training time, we apply permutations to anonymized identities to prevent the model from relying on consistent identities. Specifically, we randomly shuffle the mapping from player names to the anonymized player identities in utterances for every iteration. For example, [“David”, “Alice”, “Thomas”] [“Player1”, “Player2”, “Player3”] [“Player3”, “Player1”, “Player2”]. This mapping permutation from the text domain is also applied to the visual position encoding and ground truth label to ensure that language and visual cues are consistently aligned. This player permutation learning forces the model to learn more generalizable representations of player interactions that do not depend on specific identifiers during the training time.
| Method | Densely Aligned? | Pronoun Coreference Resolution (%) | |
| YouTube | Ego4D | ||
| BERT [24] | - | 60.3 | 47.3 |
| BERT + DINOv2 [38] | ✗ | 58.2 | 46.4 |
| BERT + MViT (Lai et al. [28]) | ✗ | 59.8 | 46.4 |
[…]
| RoBERTa [33] | - | 69.0 | 48.2 |
| RoBERTa + DINOv2 [38] | ✗ | 68.6 | 46.4 |
| RoBERTa + MViT (Lai et al. [28]) | ✗ | 69.5 | 49.1 |
[…]
| ELECTRA [13] | - | 62.5 | 44.6 |
| ELECTRA + DINOv2 [38] | ✗ | 61.1 | 42.9 |
| ELECTRA + MViT (Lai et al. [28]) | ✗ | 62.1 | 43.8 |
[…]
5 Experiments
5.1 Implementation
We adopt the language model as pre-trained BERT [24], RoBERTa [33], and ELECTRA [13] which are based on masked-language modeling. The proposed model is trained by Adam optimizer [25] with a learning rate of 5e-6 for the language model and 5e-5 for the other parts. We use a batch size of 16. We leverage about 3 seconds of video frames (frame interval 0.4s) that correspond to the timeline of the utterance. We use the preceding and following 5 utterances for encoding conversation context. The detailed network structures are described in the supplementary material.
5.2 Performance Comparison
We measure the identity classification accuracies for our curated tasks: speaking target identification, pronoun coreference resolution, and mentioned player prediction.
Table 2 shows the experimental results for speaking target identification on YouTube and Ego4D datasets with different language models. We compare our proposed baselines with the recent multimodal model [28] (i.e., Language Model + MViT [15]) for social deduction games. In addition, we further adopt DINOv2 [38] which is a powerful versatile visual feature generally used for various downstream tasks. Note that both comparison methods cannot leverage densely aligned language-visual representations. As shown in the table, these methods are not effective in improving upon the language models alone. This reflects that they are not able to figure out who the speaker is and who their gestures are directed at, in correspondence with the language domain. In contrast, our baselines leveraging densely aligned multimodal representations consistently enhance the language models for this task.
Table 3 and 4 show the performance comparison results for the pronoun coreference resolution and mentioned player prediction, respectively. As in the previous task, we conduct experiments with three different language models on two datasets. Competing methods do not perform effectively against the language baselines, but the proposed multimodal baseline performs consistently better than language baselines and the other methods. Our multimodal approach demonstrates the effectiveness of non-verbal visual cues in addressing our introduced social tasks.
| Method | Densely Aligned? | Mentioned Player Prediction (%) | |
| YouTube | Ego4D | ||
| BERT [24] | - | 54.6 | 46.2 |
| BERT + DINOv2 [38] | ✗ | 54.4 | 47.4 |
| BERT + MViT (Lai et al. [28]) | ✗ | 53.1 | 46.2 |
[…]
| RoBERTa [33] | - | 59.9 | 50.0 |
| RoBERTa + DINOv2 [38] | ✗ | 60.7 | 50.0 |
| RoBERTa + MViT (Lai et al. [28]) | ✗ | 60.6 | 51.3 |
[…]
| ELECTRA [13] | - | 55.7 | 42.3 |
| ELECTRA + DINOv2 [38] | ✗ | 56.1 | 43.6 |
| ELECTRA + MViT (Lai et al. [28]) | ✗ | 55.3 | 41.0 |
[…]
| Target Task | Method | Accuracy (%) |
| Speaking Target Identification | w/o Visual Features | 65.8 |
| w/o Gesture Feature | 69.6 | |
| w/o Gaze Feature | 70.2 | |
[…]
| Pronoun Coreference Resolution | w/o Visual Features | 60.3 |
| w/o Gesture Feature | 64.9 | |
| w/o Gaze Feature | 66.5 | |
[…]
| Mentioned Player Prediction | w/o Visual Features | 54.6 |
| w/o Gesture Feature | 55.8 | |
| w/o Gaze Feature | 56.2 | |
[…]
| Target Task | Conversation Context | Accuracy (%) | |
| Preceding | Following | ||
| Speaking Target Identification | ✗ | ✗ | 40.2 |
| ✓ | ✗ | 59.1 | |
[…]
| Pronoun Coreference Resolution | ✗ | ✗ | 51.1 |
| ✓ | ✗ | 63.4 | |
[…]
| Mentioned Player Prediction | ✗ | ✗ | 36.2 |
| ✓ | ✗ | 47.3 | |
[…]
5.3 Effects of Visual Features
We conduct ablation studies on visual feature types to analyze the contribution of each component in our baseline model. Table 5 shows the performance results according to the types of encoded non-verbal cues (i.e., gesture and gaze features) for our social tasks. In these experiments, we adopt the BERT-based baseline for evaluation on YouTube dataset. As shown in the table, the proposed baseline using both gesture and gaze features generally achieves good results. In the case of pronoun coreference resolution, the model utilizing only a gesture feature (i.e., model w/o gaze feature) achieves slightly better performance than our final baseline model. We hypothesize this is because players frequently use hand gestures while referring to other players in the group social deduction game setting.
5.4 Effects of Conversation Context
Conversational context is typically helpful in understanding a single utterance. To analyze the effects of conversation context for our tasks, we conduct ablation experiments for preceding and following contexts of the target utterance. Table 6 shows the results of different contexts using the BERT-based our baseline on YouTube dataset. A model “w/o preceding and following contexts” is one that uses only the target utterance in terms of language. The last model “w/ both preceding and following contexts” is our proposed baseline. As shown in the table, leveraging both contexts shows the best results for all social tasks. It is noteworthy that the advantage of using the following context is relatively small in the pronoun coreference resolution task compared to the other tasks. It is because the nature of this task is to resolve the reference of pronouns against people that usually appear in the preceding context.
| Target Task | Player Permutation Learning | Accuracy (%) |
| Speaking Target Identification | ✗ | 70.8 |
[…]
| Pronoun Coreference Resolution | ✗ | 51.5 |
[…]
| Mentioned Player Prediction | ✗ | 52.7 |
[…]
5.5 Effects of Permutation Learning
To validate the effectiveness of our player permutation learning which shuffles anonymized player identities, we conduct ablation experiments by training models with and without permutation. Table 7 shows the experiment results for three tasks with our BERT-based baseline on YouTube dataset. As shown in the table, the permutation learning approach consistently improves the performances for all tasks, implying it helps the model to learn more generalizable representations of player interactions. Note that we apply this player permutation learning for all comparison methods in Tables 2, 3, and 4 for fair performance comparisons.
5.6 Qualitative Results
Figure 4 shows examples of three social tasks and their qualitative results according to the use of visual cues. We utilize BERT as the language model for this experiment. As shown in the figure, our multimodal baseline leveraging both language and visual cues in a dense alignment manner can correct the inference when the language model alone fails. The qualitative results show that visual features aligned to utterances provide complementary information to disambiguate referents in social interactions.
6 Conclusion
We introduce three new challenging tasks in social deduction games: speaking target identification, pronoun coreference resolution, and mentioned player prediction - all of which require understanding the fine-grained verbal and non-verbal dynamics between multiple people. We curate extensive dataset annotations for our new social tasks and further propose a novel multimodal baseline that establishes dense language-visual alignments between spoken utterances and player visual features. This approach enables modeling multi-party social interactions through verbal and non-verbal communication channels simultaneously. Experiments show consistent and considerable performance improvements of our multimodal baselines over other approaches without both modalities, and without dense multimodal alignment. Furthermore, extensive ablation studies are conducted to validate the effectiveness of our baseline components. We will publicly release the benchmarks and code to facilitate further research in this direction.